OpenStudy (anonymous):

let f(x)=3x-4 and g(x)=-x^2 evaluate the composite function (g*g)(2)

OpenStudy (anonymous):

Do you mean (g o g)(x)?

OpenStudy (anonymous):

uh yeah

OpenStudy (anonymous):

It's the same as g( g(x) )

OpenStudy (anonymous):

except the x says two

OpenStudy (anonymous):

Yup, x is just a placeholder for any value.

OpenStudy (anonymous):

i tried that but it gave me an answer that wasnt an option.

OpenStudy (anonymous):

You tried \[(x ^{2})^{2}\] with x=2?

OpenStudy (anonymous):

nope one sec

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem. Here's a link that explains what's going on. One function basically becomes the input to a different function. http://www.mathsisfun.com/sets/functions-composition.html

OpenStudy (anonymous):

what if its (g o f)(0)

OpenStudy (anonymous):

Did you read the link?

OpenStudy (anonymous):

yeah but its still really confusing

OpenStudy (anonymous):

(g o f)(x) becomes g( f(x) ). That means f(x) is the input for the function g. Once you have it in this form, just plug in zero for x.

OpenStudy (anonymous):

If g(x) = 5x, then g(2)=5*2. And g(x+2)=5(x+2). Whatever is the input for a function just replaces the independent variables in the function.

OpenStudy (anonymous):

so then if f(x)=3x-4 and g(x)=-x^2 then i would put -x^2(3x-4) and then input 0 for my x's?

OpenStudy (anonymous):

in the end wouldnt the answer end up being 0?

OpenStudy (anonymous):

No, in that you are simply multiply the expression together. You need to make f(x), which equals 3x-4, the input for g(x). You replace any x's in g(x) with whatever f(x) is, in this case 3x-4.

OpenStudy (anonymous):

g(x) = x^2. We know that g(2) = (2)^2. And we know that g( x+2) = (x+2)^2. Do you see what is happening?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

\[g(x)=x ^{2}\]\[g(2)=2^{2}\]\[g(x+2)=(x+2)^{2}\] \[f(x)=3x-4\]\[g(f(x))=(f(x))^{2}\]\[g(3x-4)=?\]

OpenStudy (anonymous):

3gx−4g? im not sure. and im sorry for asking all the questions. the system didnt teach me all this they only taut me the simple stuff

OpenStudy (anonymous):

It's okay. I appreciate those willing to ask questions instead of just saying, "I don't get it." Do you see how every time before we took what was in the parenthesis and squared it? That is what g(x) is, x^2. So what we're doing now is the same thing; we take what's in the parenthesis and squaring it, because that is what g(x) is. So: \[g(x)=x^{2}\]\[g(f(x))=g(3x-4)=(?)^{2}\]

OpenStudy (anonymous):

i got g(3x-4)

OpenStudy (anonymous):

\[g(x)=x^{2}\]\[g(f(x))=g(3x-4)=(3x-4)^{2}\]